Can someone please help me with this question please and thanku
Answeres with steps with highly be appreciated
When 2x^3+x^2-2kx+f is divided by x-1, the remainder is -4 and when it is divided by x+2 the remainder is 11 determine the values of k and f

Answer by stanbon(75887)

When 2x^3+x^2-2kx+f is divided by x-1, the remainder is -4 and when it is divided by x+2 the remainder is 11 determine the values of k and f
-----------------
Using synthetic division::
1)....2....1....-2k....f
......2....3.....-2k+3..|..-2k+3+f
-------
-2)....2....1.....-2k.....f
.......2....-3....-2k+6..|..4k-12+f
--------
Equations:
-2k+3 + f = -4
4k-12+f = 11
-------------------
Rearrange:
-2k + f = -7
4k + f = 23
-----------------
Subtract and solve for "k"::
6k = 30
k = 5
-----
Solve for f""
4k + f = 23
20 + f = 23
f = 3
-------------
Cheers,
Stan H.
---------------